Tohban Report 2012-01-18

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 11 Jan 2012
End Date: 18 Jan 2012
Tohban: Hazel Bain
Tohban email:
Next Tohban: Claire Raftery
List all reports


Solar Activity

Solar activity this week was low to medium. There are currently 12 AR's on disk, although 5 or 6 of these will rotate over the limb within the next 24 hours. The most active region will remain on disk for another week. RHESSI unfortunately did not catch either of the 2 M class flares, the second of which occurred within hours of beginning the RHESSI anneal process on the 16th Jan 2012.

How many GOES flares occurred?

Flares above B, C, M, X class were      7    26     2     0

And how many of these are listed in the RHESSI flare list?

Flares above B, C, M, X class were      3    15     0     0

And how many had EXCELLENT coverage?

Flares above B, C, M, X class were      2     7     0     0

There were RHESSI flares/GOES flares 198 / 35 over the time range 11-Jan-12 18-Jan-12


The SSR has a max of around 15%. During the anneal an increase in SAS data will be acquired.

Data Gaps

There were a number of 300 s data gaps on the 12th Jan 2012.

DATA GAPS FOR 2012/01/12 TIME RANGE: 2012-01-12T00:00:00.000 -- 2012-01-13T00:00:00.000 GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 13 GAP START TIME GAP END TIME GAP (SEC) 2012-01-12T13:30:00.000 -- 2012-01-12T13:35:00.000 300.00000 2012-01-12T13:55:00.000 -- 2012-01-12T14:00:00.000 300.00000 2012-01-12T14:15:00.000 -- 2012-01-12T14:20:00.000 300.00000 2012-01-12T14:35:00.000 -- 2012-01-12T14:40:00.000 300.00000 2012-01-12T16:00:00.000 -- 2012-01-12T16:05:00.000 300.00000 2012-01-12T17:50:00.000 -- 2012-01-12T17:55:00.000 300.00000 2012-01-12T18:00:00.000 -- 2012-01-12T18:05:00.000 300.00000 2012-01-12T18:25:00.000 -- 2012-01-12T18:30:00.000 300.00000 2012-01-12T18:35:00.000 -- 2012-01-12T18:40:00.000 300.00000 2012-01-12T18:55:00.000 -- 2012-01-12T19:05:00.000 600.00000 2012-01-12T21:30:00.000 -- 2012-01-12T21:35:00.000 300.00000 2012-01-12T23:05:00.000 -- 2012-01-12T23:10:00.000 300.00000 2012-01-12T23:40:00.000 -- 2012-01-12T23:45:00.000 300.00000 GAPS IN APP_ID = 154 (VC1-PMTRAS) WITH PACKET RATE LT 4 N_GAPS 1 GAP START TIME GAP END TIME GAP (SEC) 2012-01-12T21:00:00.000 -- 2012-01-12T21:05:00.000 300.00000 NO GAPS IN APP_ID = 102 (VC3-MONITOR RATES)

Detector issues

In the 24 hours prior to the start of the anneal, detector 8 was showing significant periods of noise similar to that seen over the previous few months.

Annealing plans

The RHESSI anneal began on the 17th Jan 2012 02:52:45 UT. The detectors were shut off and the cryocooler power was reduced to 50W and then 20W. On the third BGS pass of the night, the coldplate heater was turned on with a set point of 50 C. On the fourth BGS contact the cryocooler was found to be turned off due to hardware Over Current trip (at 2012-017-06:32:36 UTC). The coldplate heater was turned off to make ensure the cold tip would not get too warm. The back orbit data gave no indication that anything unusual happened to the cryocooler leading up to the trip. The decision was made to leave the cryocooler off but to leave the coldplate heater on, with a low set point of 10 C on, to ensure the detectors would not settle at a temperature around 150 - 170 K, which is a very bad temperature for radiation damage. Currently the cold plate and cold tip temperatures are as follows:

Cold Tip Temperature 1 253.918 K Cold Tip Temperature 2 253.044 K Cold Plate Temperature 1 274.251 K Cold Plate Temperature 2 272.837 K

Following a discussion with David Curtis and others it was decided to continue turn the cryocooler back on and continue with the anneal during the first BGS pass this evening (i.e. 2012/019 02:29:17 UT).

Personal tools